G NEN (a) RELATE T oA GEOmufn SERIES OFT FORM no VRLUES WHAT SERIES 5700 んJo CONVERGF (C) um? WHAT
Answer #1

() SOlubon :a) Gven CO (z2+1) + O + + (z+(z2+) cahere TtZ ho For the seies Conveene (b) T</1tZI Z+1>1 -1 z20 OJ z2+2 O E(-0 0Ce) Foran -1 1- h= z2+1 - 1 *-Jt2 - L 2 z z2 11 2. Awer h=1 -IN 1